OpenStudy (anonymous):

test for symmetry with respect to each axis and to the origin. y=x^3+x

OpenStudy (anonymous):

first I should see if its symmetric with the origin..i believe you change the x and y values signs.. can you show that first?

OpenStudy (mathteacher1729):

There are two ways to do this -- visually (using a grapher) or algebraically. Are you supposed to do one or the other?

OpenStudy (anonymous):

I am supposed to show my work, so yeah algebraically

OpenStudy (anonymous):

\[(-y) = (-x)^3 + (-x) \\ -y = -x - x\] Compare that with what you get originally

OpenStudy (anonymous):

So far, to check if symmetric on origin, I have y=x^3+x, then I made it -y=(-x)^3 - x

OpenStudy (anonymous):

that doesn't match, so it isn't symmetric with the origin?

OpenStudy (mathteacher1729):

Ok, that makes sense. Are you allowed to verify your work or use a grapher as you are solving the problem? (It will help you make sure your answers are correct.) There are two main tests: 1) Symmetric about origin Everywhere you see an x, replace it with a y and vice versa. 2) Symmetric about the y-axis Everywhere you see an x , replace it with -x 3) If the function was symmetric about the y axis... it wouldn't be a function (wouldn't be a function of x anyway). (why is this?)

OpenStudy (anonymous):

I have to show it on paper. I already checked origin, and found it isn't symmetric about it,right? Now can you help me find if it is symmetric with x and y? :)

OpenStudy (anonymous):

So if symmetric with y axis? would I have: y= -x - x after simplification?

OpenStudy (mathteacher1729):

@schmidtdancer yup, you showed it's not symmetric about the origin. My first step was actually finding if it was symmetric about the line y = x.

OpenStudy (mathteacher1729):

y = x^3 + x replace x with -x y = (-x)^3 + (-x) y = -x -x y = -2x Clearly this is not the same.

OpenStudy (anonymous):

ok, thats what I had. so not symmetric about origin OR y-axis. Now, lastly, do i check if symmetric about x-axis?

OpenStudy (mathteacher1729):

By the way there are LOTS and LOTS of worked examples here: http://tutorial.math.lamar.edu/Classes/Alg/Symmetry.aspx They might be useful to reference against more complex problems.

OpenStudy (anonymous):

ok.

OpenStudy (mathteacher1729):

y = (-x)^3 + (-x) y = (-1)^3(x)^3 - x notice that we are raising -x to the third. This is just negative one to the third times x to the third. y = -x^3 -x I think it might be too early for me to be on open study. I feel I'm making too many errors. :(

OpenStudy (anonymous):

So we are wrong??

OpenStudy (anonymous):

Oh that was for symmetry on y-axis.... But it is still no? right,

OpenStudy (mathteacher1729):

Here's how it breaks down. 1) Symmetric about the origin = YES 2) Symmetric about x axis = NO 3) Symmetric about y axis = NO I'm going to write up the reasons why and double-check them. I apologize for any confusion. :( I'm usually more on the ball but I just woke up.

OpenStudy (anonymous):

No prob! Are u sure about those?

OpenStudy (mathteacher1729):

Hey sorry I'm back.

OpenStudy (mathteacher1729):

Symmetric about origin? replace x with -x replace y with -y If the equation is the same, then YES. Otherwise, NO. 1) Origin y = x^3 + x (-y) = (-x)^3 + (-x) replacement of variables -y = -x^3 - x note that a negative raised to an ODD power is always negative. y = x^3 +x dividing through by -1 gives us the original equation. YES 2) x-axis replace y with -y If the equation is the same, then YES. Otherwise, NO. y = x^3 + x -y = x^3 + x replacement of variable the equations are NOT the same, you can't get rid of the negative in front of the y. NO 3) y-axis replace x with -x y = x^3 + x y = (-x)^3 + (-x) replacement of variable y = -x^3 - x note that a negative raised to an ODD power is always negative. the equations are NOT the same, you can't get rid of the negative in front of the xs. NO

OpenStudy (mathteacher1729):

I usually don't like to type up full replies like that with such explicit answers, but in this case I felt it was needed because I was being confusing in the beginning. Hope this helps, and please check out http://tutorial.math.lamar.edu/Classes/Alg/Symmetry.aspx for even more examples. :)